Mountain Top Markets has total assets of $48,700, net working capital of $1,100, and retained earnings of $21,200. The firm has 12,500 shares of stock outstanding with a par value of $1 per share and a market value of $7.10 per share. The stock was originally issued to the firm's founders at par value. What is the market-to-book ratio?

A) 3.19

B) 2.22

C) 2.78

D) 3.03

E) 2.63
Answer #1

Answer is 2.63

market-to-book ratio=Market price per share / Book Value

Book Value per share=networth/No of Share outstanding

Book Value per share=(12500*1$+$21200)/12500=$2.70

market-to-book ratio=7.10/2.70=2.63